Gasoline prices are rising at a rate of 0.5% per month. What is their doubling time? By what factor will prices increase in 3 years? In 5 years?
I think the doubling time is 138.98 months? and the increases are 1.5% and 2.5%?

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

P=Po(1+.005)^n
P/Po=2
2=1.005^n
ln2=n*ln(1.005)
n=138.98
for 1.5%
P=Po(1.015)^n
ln2=n*ln(1.015)
n=46.56 months
for 2.5%
ln2=n*ln(1.025)
n=28.07 months.
3 years for 0.5%
1.005^36=1.20 times the original
5 years for 0.5%
1.005^60=1.35 times the original

Answer by MathTherapy(10556)


Gasoline prices are rising at a rate of 0.5% per month. What is their doubling time? By what factor will prices increase in 3 years? In 5 years?
I think the doubling time is 138.98 months? and the increases are 1.5% and 2.5%?
You're correct! The doubling time is 138.98 months.
However, for 3 and 5 years, the price will increase by factors of: